OpenStudy (anonymous):

If f −1(x) = x2 + 5, then what is the value of function f(x)?

OpenStudy (anonymous):

please answer quickly

Parth (parthkohli):

I'd answer you after 30 minutes. :-)

OpenStudy (anonymous):

the exam is a test of your knowledge, not a test of the knowledge you can gather during the exam.

OpenStudy (anonymous):

@ParthKohli wait an hour. or an hour and a half.

OpenStudy (anonymous):

ok

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (phi):

\[ f ^{−1}(x) = x^2 + 5 \] one way is write it as y= x^2 +5 now *swap x and y* x= y^2 +5 last step: solve for y. that will be your function